In a circle with radius 13, two parallel chords with length 10 are joined to create a rectangle. Find the area of this rectangle.

Answer by richwmiller(17219)

one side is 10 the diagonal is 26
26^2=10^2+b^2
676=100+b^2
576=b^2
24=b
10*24=240 square units is the area